OpenStudy (anonymous):

Sketch the solution to the initial value problem and determine its maximum value. see post below for equation

OpenStudy (anonymous):

\[dy/dt=2y-2yt \]

OpenStudy (anonymous):

where y(0)=3

OpenStudy (myname):

First, you have to find y(t)

OpenStudy (myname):

\[dy/dt = 2y-2yt\] Now, you separate the y's and t's \[dy/dt=2y(1-t)\] Next, divide both sides by 2y and multiply both sides by dt \[dy/2y=(1-t)dt\] Now we need to integrate both sides. \[\int\limits_{}^{}dy/2y=\int\limits_{}^{}(1-t)dt\] \[\ln(y)/2=t-t^2/2 +c\] \[\ln(y)=2t-t^2+2c\] e^(lny)=e^(2t-t^2 +2c) [e^lny = y] So, y= e^(2t-t^2+2c) ----- By using laws of exponent--- y = (e^2t/e^t^2) * e^2c As y(0)=3 or f(0)=3 3 = (e^0/e^0)*e^2c 3=e^2c c=3 -----------------------> as e^2c is just a constant. Thus, y = e^(2t-t^2+6)

OpenStudy (myname):

In order to find the maximum first you find the critical points[points where the first derivative is 0.]

OpenStudy (myname):

Hint : -find the first derivative -solve for t as y(prime) equals 0. And that will be your critical points. -plug in the critical points in the y function, then you will get the maximum.(the maximum value obtained is going to be the max.)